# Future Impact Decomposition in Request-level Recommendations

### v4

@@ Appendix D The Details of all Methods
All models and baselines use the same user request encoder and the main difference locates in the design of the learning framework and the actor/critic design. We list details of these differences as follows:
@@
Appendix E Full training curves
In Figure 8 we present the training curves of all metrics (reward, variance, and depths) and training losses (critic loss and actor loss) in the online learning at both user environments. We include HAC as the strongest baseline and compare our ItemA2C alternatives with it. We can see that ItemA2C-M outperforms others on all metrics in the ML1M environment and significantly outperforms others in KuaiRand. Noted that $\mathcal{L}_{\mathrm{actor}}$ and $\mathcal{L}_{\mathrm{critic}}$ of gradually converge during the training process. We can also observe the lower reward variance of ItemA2C-M’s across users, indicating that the learned policy is more capable of providing good and stable actions under different user states.

### v5

@@ Appendix C The Details of all Methods
All models and baselines use the same user request encoder and the main difference locates in the design of the learning framework and the actor/critic design. Except for the SlateQ that generate K items simultaneously and optimizes the joint action, all other baselines output a single vector action and adopt KNN-style top-K selection. We list details of these differences as follows:
@@
Appendix D Full training curves
In Figure 8 we present the training curves of all metrics (reward, variance, and depths) and training losses (critic and actor) in the online learning at both user environments. We include HAC as the strongest baseline and compare our ItemA2C alternatives with it. We can see that ItemA2C-M outperforms others on all metrics in the ML1M environment and significantly outperforms others in KuaiRand. Noted that $\mathcal{L}_{\mathrm{actor}}$ and $\mathcal{L}_{\mathrm{critic}}$ of gradually converge during the training process. We can also observe the lower reward variance of ItemA2C-M’s across users, indicating that the learned policy is more capable of providing good and stable actions under different user states.